OpenStudy (anonymous):

solve for x: (x+4)/3=2

OpenStudy (mathmale):

You've done a nice job presenting this problem in that you've enclosed x+4 inside parentheses. This is the proper thing to do. To make this solution easier, why not multiply both sides of the given equation by 3, to remove the fractions? Type out your result here, please.

OpenStudy (akashdeepdeb):

\[\frac{x+4}{3} = 2\] Multiply both sides by 3 [To get rid of the '3' on the LHS] You'll get: \[\frac{(x+4)*3}{3} = 2*3\] Which further equals: \[x+4 = 6\] Can you complete this now, using rules of simultaneous equations? :)

OpenStudy (mathmale):

Mr. Man: Ask yourself: What is the goal of this problem? How are you going to solve the equation x + 4 = 6?

OpenStudy (anonymous):

x=6-4=2